OpenStudy (anonymous):

Solve the equation using the quadratic formula. 2x^2 = -12x - 3

OpenStudy (anonymous):

arrange the equation 2x^2 = -12x - 3 to 2x^2 +12x + 3=0 where a=2, b=12 and c=3 ax^2 +bx + c=0 use the quadratic formula \[x=\frac{ -b -+\sqrt{b ^{2}-4ac}}{ 2a }\]